OpenStudy (anonymous):

Evaluate and write your answer in scientific notation:(4 x 10^2)(15 x 10^3) A. 60 x 10^6 B. 6 x 10^6 C. 6 x 10^5

OpenStudy (anonymous):

(4 x 10^2) is 4 x 100 since 10^2=100 \[\therefore\](4 x 10^2)=400 (15 x 10^3) is 15 x 1000 since 10^3=1000 \[\therefore\] (15 x 10^3)=15000 Now since you have two brackets next to each other you multiply them. So you have 400 x 15000 Since 400 x 15000=6000000 you can simplify that by writing 6x10^6 This means that you have a 6 followed by 6 zeros. This is because you're multiplying 10^6 which is 1000000 by 6.